HPF Workshop: Final Program

Chilworth Manor, Southampton, UK.

11 October 1995

09.30 - 10.00: REGISTRATION

10.00 - 11.00: Welcome + keynote address

Welcome address Tony Hey (Head of Department, Electronics and Comupter Science, U. Southampton),

Vienna Fortran extensions for HPF2 Barbara Chapman (Director, European Centre for Parallel Computing at Vienna, Austria)

11.00 - 11.30: BREAK

11.30 - 13.00: HPF systems

shpf: a Subset HPF compilation system John Merlin and Bryan Carpenter (U. Southampton)

Optimisations in the HPF compilation system ADAPTOR Thomas Brandes (GMD, Germany)

The N.A. Software HPF Toolset Mike Delves (N.A. Software, Liverpool)

13.00 - 14.00: LUNCH

14.00 - 15.30: HPF user experiences

Parallel Integrated Forecasting System (IFS): Message-passing versus HPF Fritz Wollenweber (European Centre for Medium-Range Weather Forecasting, Reading)

Practical experiences in HPF David Greco (CRS4, Cagliari, Italy)

Porting the GENESIS benchmarks to HPF.

Vladimir Getov (U. Southampton + U. Westminster) and Bryan Carpenter (U. Southampton)

15.30 - 16.00: BREAK

16.00 - 17.30:

HPC centres experience with HPF Adam Marshal (University of Liverpool) Alistair Ewing (Edinburgh Parallel Computing Centre)

Panel discussion 17.30 :CLOSE OF WORKSHOP

2